Multiply 2m(5m-3)(4m+7)
Answer by fractalier(2101)

Let's do it in stages...
2m(5m-3)(4m+7) =
2m(20m^2 + 35m - 12m - 21) =
2m(20m^2 + 23m - 21) =
40m^3 + 46m^2 - 42m